# Efficiently Parameterized Neural Metriplectic Systems

### v2

@@ 1 Introduction
The theory of metriplectic, also called GENERIC, systems [20, 4] provides a principled formalism for encoding dissipative dynamics in terms of complete thermodynamical systems that conserve energy and produce entropy. By formally expressing the reversible and irreversible parts of state evolution with separate algebraic brackets, the metriplectic formalism provides a general mechanism for maintaining essential conservation laws while simultaneously respecting dissipative effects. Thermodynamic completeness implies that any dissipation is caught within a metriplectic system through the generation of entropy, allowing for a holistic treatment which has already found use in modeling fluids [19, 18], plasmas [11, 17], and kinetic theories [10, 9].
@@ 2 Previous and Related Work
Attempts to learn metriplectic systems using soft constraints rely on relaxing the degeneracy conditions ${}L\nabla S=~{}M\nabla E=~{}0$. This is the approach taken in [8], termed SPNN, which learns an almost-metriplectic model parameterized with generic neural networks through a simple $L^{2}$ penalty term in the training loss, $\mathcal{L}_{\mathrm{pen}}=\mathopen{}\mathclose{{}\left|~{}L\nabla E}\right|^{2}+\mathopen{}\mathclose{{}\left|~{}M\nabla S}\right|^{2}.$ This widens the space of allowable network parameterizations for the approximate operators ${}L,~{}M$. While the resulting model violates the first and second laws of thermodynamics, the authors show that reasonable trajectories are still obtained, at least when applied within the range of timescales used for training. A similar approach is taken in [7], which targets larger problems and develops an almost-metriplectic model reduction strategy based on the same core idea.
@@
Perhaps the first example of learning metriplectic systems from data was given in [3] in the context of system identification. Here, training data is assumed to come from a finite element simulation, so that the discrete gradients of energy and entropy can be approximated as ${}\nabla E(~{}x)=~{}A~{}x,\nabla S(~{}x)=~{}B~{}x$. Assuming a fixed form for ${}L$ produces a constrained learning problem for the constant matrices ${}M,~{}A,~{}B$ which is solved to yield a provably metriplectic surrogate model. Similarly, the work [24] learns ${}M,E$ given ${}L,S$ by considering a fixed block-wise decoupled form which trivializes the degeneracy conditions, i.e. ${}L=[\star\,\,~{}0;~{}0\,\,~{}0]$ and ${}M=[~{}0\,\,~{}0;~{}0\,\,\star]$. This line of thought is continued in [30] and [31], both of which learn metriplectic systems with neural network parameterizations by assuming this decoupled block structure. A somewhat broader class of metriplectic systems are considered in [6] using tools from exterior calculus, with the goal of learning metriplectic dynamics on graph data. This leads to a structure-preserving network surrogate which scales linearly in the size of the graph domain, but also cannot express arbitrary metriplectic dynamics due to the specific choices of model form for ${}L,~{}M$.
@@
A particularly inspirational method for learning general metriplectic systems was given in [13] and termed GNODE, building on parameterizations of metriplectic operators developed in [21]. GNODE parameterizes learnable reversible and irreversible bracket generating matrices ${}L,~{}M$ in terms of state-independent tensors ${}\xi\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)^{\otimes 3}$ and ${}\zeta\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)^{\otimes 4}$: for $1\leq\alpha,\beta,\gamma,\mu,\nu\leq n$, the authors choose $L_{\alpha\beta}(~{}x)=\sum_{\gamma}\xi_{\alpha\beta\gamma}\partial^{\gamma}S$ and $M_{\alpha\beta}(~{}x)=\sum_{\mu,\nu}\zeta_{\alpha\mu,\beta\nu}\partial^{\mu}E\partial^{\nu}E,$ where $\partial^{\alpha}F=\partial F/\partial x_{\alpha}$, ${}\xi$ is totally antisymmetric, and ${}\zeta$ is symmetric between the pairs $(\alpha,\mu)$ and $(\beta,\nu)$ but antisymmetric within each of these pairs. The key idea here is to exchange the problem of enforcing degeneracy conditions ${}L\nabla E=~{}M\nabla S=~{}0$ in matrix fields ${}L,~{}M$ with the problem of enforcing symmetry conditions in tensor fields ${}\xi,~{}\zeta$, which is comparatively easier but comes at the expense of underdetermining the problem. In GNODE, enforcement of these symmetries is handled by a generic learnable 3-tensor $\tilde{~{}\xi}\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)^{\otimes 3}$ along with learnable matrices ${}D\in\mathrm{Sym}_{r}\mathopen{}\mathclose{{}\left(\mathbb{R}}\right)$ and ${}\Lambda^{s}\in\mathrm{Skew}_{n}\mathopen{}\mathclose{{}\left(\mathbb{R}}\right)$ for $1\leq s\leq r\leq n$, leading to the final parameterizations $\xi_{\alpha\beta\gamma}=\frac{1}{3!}\mathopen{}\mathclose{{}\left(\tilde{\xi}_{\alpha\beta\gamma}-\tilde{\xi}_{\alpha\gamma\beta}+\tilde{\xi}_{\beta\gamma\alpha}-\tilde{\xi}_{\beta\alpha\gamma}+\tilde{\xi}_{\gamma\alpha\beta}-\tilde{\xi}_{\gamma\beta\alpha}}\right)$ and $\zeta_{\alpha\mu,\beta\nu}=\sum_{s,t}\Lambda_{\alpha\mu}^{s}D_{st}\Lambda_{\beta\nu}^{t}.$ Along with learnable energy and entropy functions $E,S$ parameterized by multi-layer perceptrons (MLPs), the data ${}L,~{}M$ learned by GNODE guarantees metriplectic structure in the surrogate model and leads to successful learning of metriplectic systems in some simple cases of interest. However, note that this is a highly redundant parameterization requiring (${n\choose 3}+r{n\choose 2}+{{r+1}\choose 2}+2$ learnable scalar functions, which exhibits $\mathcal{O}(n^{3}+rn^{2})$ scaling in the problem size because of the necessity to compute and store (${n\choose 3}$ entries of ${}\xi$ and $r{n\choose 2}$ entries of ${}\Lambda$. Additionally, the assumption of state-independence in the bracket generating tensors ${}\xi,~{}\zeta$ is somewhat restrictive, limiting the class of problems to which GNODE can be applied.
@@
A related approach to learning metriplectic dynamics with hard constraints was seen in [32], which proposed a series of GFINN architectures depending on how much of the information ${}L,~{}M,E,S$ is assumed to be known. In the case that ${}L,~{}M$ are known, the GFINN energy and entropy are parameterized with scalar-valued functions $f\circ~{}P_{\mathrm{ker}~{}A}$ where $f:\mathbb{R}^{n}\to\mathbb{R}$ ($E$ or $S$) is learnable and ${}P_{\mathrm{ker}~{}A}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is orthogonal projection onto the kernel of the (known) operator ${}A$ (${}L$ or ${}M$). It follows that the gradient $\nabla\mathopen{}\mathclose{{}\left(f\circ~{}P_{\mathrm{ker}~{}A}}\right)=~{}P_{\mathrm{ker}~{}A}\nabla f\mathopen{}\mathclose{{}\left(P_{\mathrm{ker}~{}A}}\right)$ lies in the kernel of ${}A$, so that the degeneracy conditions are guaranteed at the expense of constraining the model class of potential energies/entropies. Alternatively, in the case that all of ${}L,~{}M,E,S$ are unknown, GFINNs use learnable scalar functions $f$ for $E,S$ parameterized by MLPs as well as two matrix fields ${}Q^{E},~{}Q^{S}\in\mathbb{R}^{r\times n}$ with rows given by ${}q^{f}_{s}=\mathopen{}\mathclose{{}\left(~{}S^{f}_{s}\nabla f}\right)^{\intercal}$ for learnable skew-symmetric matrices ${}S^{f}_{s}$, $1\leq s\leq r$, $f=E,S$. Along with two triangular $(r\times r)$ matrix fields ${}T_{~{}L},~{}T_{~{}M}$, this yields the parameterizations ${}L(~{}x)=~{}Q^{S}(~{}x)^{\intercal}\mathopen{}\mathclose{{}\left(~{}T_{~{}L}(~{}x)^{\intercal}-~{}T_{~{}L}(~{}x)}\right)~{}Q^{S}(~{}x)$ and ${}M(~{}x)=~{}Q^{E}(~{}x)^{\intercal}\mathopen{}\mathclose{{}\left(~{}T_{~{}M}(~{}x)^{\intercal}~{}T_{~{}M}(~{}x)}\right)~{}Q^{E}(~{}x)$, which necessarily satisfy the symmetries and degeneracy conditions required for metriplectic structure. GFINNs are shown to both increase expressivity over the GNODE method as well as decrease redundancy, since the need for an explicit order-3 tensor field is removed and the reversible and irreversible brackets are allowed to depend explicitly on the state ${}x$. However, GFINNs still exhibit cubic scaling through the requirement of $rn(n-1)+r^{2}+2=\mathcal{O}\mathopen{}\mathclose{{}\left(rn^{2}}\right)$ learnable functions, which is well above the theoretical minimum required to express a general metriplectic system and leads to difficulties in training the resulting models.
@@
Finally, it is worth mentioning the closely related line of work involving model reduction for metriplectic systems, which began with [22]. As remarked there, preserving metriplecticity in reduced-order models (ROMs) exhibits many challenges due to its delicate requirements on the kernels of the involved operators. There are also hard and soft constrained approaches: the already mentioned [7] aims to learn a close-to-metriplectic data-driven ROM by enforcing degeneracies by penalty, while [5] directly enforces metriplectic structure in projection-based ROMs using exterior algebraic factorizations. The parameterizations of metriplectic data presented here are related to those presented in [5], although NMS does not require access to nonzero components of $\nabla E,\nabla S$.
@@ 3.1 Exterior algebra
Developing these metriplectic expressions will require some basic facts from exterior algebra, of which more details can be found in, e.g., [28, Chapter 19]. The basic objects in the exterior algebra $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}V$ over the vector space $V$ are multivectors, which are formal linear combinations of totally antisymmetric tensors on $V$. More precisely, if $I(V)$ denotes the two-sided ideal of the free tensor algebra $T(V)$ generated by elements of the form ${}v\otimes~{}v$ (${}v\in V$), then the exterior algebra is the quotient space $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}V\simeq T(V)/I(V)$ equipped with the antisymmetric wedge product operation $\wedge$. This graded algebra is equipped with natural projection operators $P^{k}:\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}V\to\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{k}V$ which map between the full exterior algebra and the $k^{\mathrm{th}}$ exterior power of $V$, denoted $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{k}V$, whose elements are homogeneous $k$-vectors. More generally, given an $n$-manifold $M$ with tangent bundle $TM$, the exterior algebra $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}\mathopen{}\mathclose{{}\left(TM}\right)$ is the algebra of multivector fields whose fiber over $x\in M$ is given by $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}T_{x}M$.
For the present purposes, it will be useful to develop a correspondence between bivectors $\mathsf{B}\in\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{2}\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)$ and skew-symmetric matrices ${}B\in\mathrm{Skew}_{n}\mathopen{}\mathclose{{}\left(\mathbb{R}}\right)$, which follows directly from Riesz representation in terms of the usual Euclidean dot product. More precisely, supposing that ${}e_{1},...,~{}e_{n}$ are the standard basis vectors for $\mathbb{R}^{n}$, any bivector $\mathsf{B}\in\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{2}T\mathbb{R}^{n}$ can be represented as $\mathsf{B}=\sum_{i<j}B^{ij}~{}e_{i}\wedge~{}e_{j}$ where $B^{ij}\in\mathbb{R}$ denote the components of $\mathsf{B}$. Define a grade-lowering action of bivectors on vectors through right contraction (see e.g. Section 3.4 of [2]), expressed for any vector ${}v$ and basis bivector ${}e_{i}\wedge~{}e_{j}$ as $\mathopen{}\mathclose{{}\left(~{}e_{i}\wedge~{}e_{j}}\right)\cdot~{}v=\mathopen{}\mathclose{{}\left(~{}e_{j}\cdot~{}v}\right)~{}e_{i}-\mathopen{}\mathclose{{}\left(~{}e_{i}\cdot~{}v}\right)~{}e_{j}.$ It follows that the action of $\mathsf{B}$ is equivalent to

### v3

@@ 1 Introduction
The theory of metriplectic, also called GENERIC, systems Morrison (1986); Grmela and Öttinger (1997) provides a principled formalism for encoding dissipative dynamics in terms of complete thermodynamical systems that conserve energy and produce entropy. By formally expressing the reversible and irreversible parts of state evolution with separate algebraic brackets, the metriplectic formalism provides a general mechanism for maintaining essential conservation laws while simultaneously respecting dissipative effects. Thermodynamic completeness implies that any dissipation is caught within a metriplectic system through the generation of entropy, allowing for a holistic treatment which has already found use in modeling fluids Morrison (1984; 2009), plasmas Kaufman and Morrison (1982); Materassi (2012), and kinetic theories Kaufman (1984); Holm et al. (2008).
@@ 2 Previous and Related Work
Attempts to learn metriplectic systems using soft constraints rely on relaxing the degeneracy conditions ${}L\nabla S=~{}M\nabla E=~{}0$. This is the approach taken in Hernández et al. (2021b), termed SPNN, which learns an almost-metriplectic model parameterized with generic neural networks through a simple $L^{2}$ penalty term in the training loss, $\mathcal{L}_{\mathrm{pen}}=\mathopen{}\mathclose{{}\left|~{}L\nabla E}\right|^{2}+\mathopen{}\mathclose{{}\left|~{}M\nabla S}\right|^{2}.$ This widens the space of allowable network parameterizations for the approximate operators ${}L,~{}M$. While the resulting model violates the first and second laws of thermodynamics, the authors show that reasonable trajectories are still obtained, at least when applied within the range of timescales used for training. A similar approach is taken in Hernández et al. (2021a), which targets larger problems and develops an almost-metriplectic model reduction strategy based on the same core idea.
@@
Perhaps the first example of learning metriplectic systems from data was given in González et al. (2019) in the context of system identification. Here, training data is assumed to come from a finite element simulation, so that the discrete gradients of energy and entropy can be approximated as ${}\nabla E(~{}x)=~{}A~{}x,\nabla S(~{}x)=~{}B~{}x$. Assuming a fixed form for ${}L$ produces a constrained learning problem for the constant matrices ${}M,~{}A,~{}B$ which is solved to yield a provably metriplectic surrogate model. Similarly, the work Ruiz et al. (2021) learns ${}M,E$ given ${}L,S$ by considering a fixed block-wise decoupled form which trivializes the degeneracy conditions, i.e. ${}L=[\star\,\,~{}0;~{}0\,\,~{}0]$ and ${}M=[~{}0\,\,~{}0;~{}0\,\,\star]$. This line of thought is continued in Xu et al. (2022) and Xu et al. (2023), both of which learn metriplectic systems with neural network parameterizations by assuming this decoupled block structure along with specialized forms for the metriplectic data. A somewhat broader class of metriplectic systems are considered in Gruber et al. (2023b) using tools from exterior calculus, with the goal of learning metriplectic dynamics on graph data. This leads to a structure-preserving network surrogate which scales linearly in the size of the graph domain, but also cannot express arbitrary metriplectic dynamics due to the specific modeling choices for ${}L,~{}M$.
@@
A particularly inspirational method for learning general metriplectic systems was given in Lee et al. (2021) and termed GNODE, building on parameterizations of metriplectic operators developed in Öttinger (2014). GNODE parameterizes learnable reversible and irreversible bracket generating matrices ${}L,~{}M$ in terms of state-independent tensors ${}\xi\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)^{\otimes 3}$ and ${}\zeta\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)^{\otimes 4}$: for $1\leq\alpha,\beta,\gamma,\mu,\nu\leq n$, the authors choose $L_{\alpha\beta}(~{}x)=\sum_{\gamma}\xi_{\alpha\beta\gamma}\partial^{\gamma}S$ and $M_{\alpha\beta}(~{}x)=\sum_{\mu,\nu}\zeta_{\alpha\mu,\beta\nu}\partial^{\mu}E\partial^{\nu}E,$ where $\partial^{\alpha}F=\partial F/\partial x_{\alpha}$, ${}\xi$ is totally antisymmetric, and ${}\zeta$ is symmetric between the pairs $(\alpha,\mu)$ and $(\beta,\nu)$ but antisymmetric within each of these pairs. The key idea here is to exchange the problem of enforcing degeneracy conditions ${}L\nabla E=~{}M\nabla S=~{}0$ in matrix fields ${}L,~{}M$ with the problem of enforcing symmetry conditions in tensor fields ${}\xi,~{}\zeta$, which is comparatively easier but comes at the expense of underdetermining the problem. In GNODE, enforcement of these symmetries is handled by a generic learnable 3-tensor $\tilde{~{}\xi}\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)^{\otimes 3}$ along with learnable matrices ${}D\in\mathrm{Sym}_{r}\mathopen{}\mathclose{{}\left(\mathbb{R}}\right)$ and ${}\Lambda^{s}\in\mathrm{Skew}_{n}\mathopen{}\mathclose{{}\left(\mathbb{R}}\right)$ for $1\leq s\leq r\leq n$, leading to the final parameterizations $\xi_{\alpha\beta\gamma}=\frac{1}{3!}\mathopen{}\mathclose{{}\left(\tilde{\xi}_{\alpha\beta\gamma}-\tilde{\xi}_{\alpha\gamma\beta}+\tilde{\xi}_{\beta\gamma\alpha}-\tilde{\xi}_{\beta\alpha\gamma}+\tilde{\xi}_{\gamma\alpha\beta}-\tilde{\xi}_{\gamma\beta\alpha}}\right)$ and $\zeta_{\alpha\mu,\beta\nu}=\sum_{s,t}\Lambda_{\alpha\mu}^{s}D_{st}\Lambda_{\beta\nu}^{t}.$ Along with learnable energy and entropy functions $E,S$ parameterized by multi-layer perceptrons (MLPs), the data ${}L,~{}M$ learned by GNODE guarantees metriplectic structure in the surrogate model and leads to successful learning of metriplectic systems in some simple cases of interest. However, note that this is a highly redundant parameterization requiring (${n\choose 3}+r{n\choose 2}+{{r+1}\choose 2}+2$ learnable scalar functions, which exhibits $\mathcal{O}(n^{3}+rn^{2})$ scaling in the problem size because of the necessity to compute and store (${n\choose 3}$ entries of ${}\xi$ and $r{n\choose 2}$ entries of ${}\Lambda$. Additionally, the assumption of state-independence in the bracket generating tensors ${}\xi,~{}\zeta$ is somewhat restrictive, limiting the class of problems to which GNODE can be applied.
@@
A related approach to learning metriplectic dynamics with hard constraints was seen in Zhang et al. (2022), which proposed a series of GFINN architectures depending on how much of the information ${}L,~{}M,E,S$ is assumed to be known. In the case that ${}L,~{}M$ are known, the GFINN energy and entropy are parameterized with scalar-valued functions $f\circ~{}P_{\mathrm{ker}~{}A}$ where $f:\mathbb{R}^{n}\to\mathbb{R}$ ($E$ or $S$) is learnable and ${}P_{\mathrm{ker}~{}A}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is orthogonal projection onto the kernel of the (known) operator ${}A$ (${}L$ or ${}M$). It follows that the gradient $\nabla\mathopen{}\mathclose{{}\left(f\circ~{}P_{\mathrm{ker}~{}A}}\right)=~{}P_{\mathrm{ker}~{}A}\nabla f\mathopen{}\mathclose{{}\left(P_{\mathrm{ker}~{}A}}\right)$ lies in the kernel of ${}A$, so that the degeneracy conditions are guaranteed at the expense of constraining the model class of potential energies/entropies. Alternatively, in the case that all of ${}L,~{}M,E,S$ are unknown, GFINNs use learnable scalar functions $f$ for $E,S$ parameterized by MLPs as well as two matrix fields ${}Q^{E},~{}Q^{S}\in\mathbb{R}^{r\times n}$ with rows given by ${}q^{f}_{s}=\mathopen{}\mathclose{{}\left(~{}S^{f}_{s}\nabla f}\right)^{\intercal}$ for learnable skew-symmetric matrices ${}S^{f}_{s}$, $1\leq s\leq r$, $f=E,S$. Along with two triangular $(r\times r)$ matrix fields ${}T_{~{}L},~{}T_{~{}M}$, this yields the parameterizations ${}L(~{}x)=~{}Q^{S}(~{}x)^{\intercal}\mathopen{}\mathclose{{}\left(~{}T_{~{}L}(~{}x)^{\intercal}-~{}T_{~{}L}(~{}x)}\right)~{}Q^{S}(~{}x)$ and ${}M(~{}x)=~{}Q^{E}(~{}x)^{\intercal}\mathopen{}\mathclose{{}\left(~{}T_{~{}M}(~{}x)^{\intercal}~{}T_{~{}M}(~{}x)}\right)~{}Q^{E}(~{}x)$, which necessarily satisfy the symmetries and degeneracy conditions required for metriplectic structure. GFINNs are shown to both increase expressivity over the GNODE method as well as decrease redundancy, since the need for an explicit order-3 tensor field is removed and the reversible and irreversible brackets are allowed to depend explicitly on the state ${}x$. However, GFINNs still exhibit cubic scaling through the requirement of $rn(n-1)+r^{2}+2=\mathcal{O}\mathopen{}\mathclose{{}\left(rn^{2}}\right)$ learnable functions, which is well above the theoretical minimum required to express a general metriplectic system and leads to difficulties in training the resulting models.
@@
Finally, it is worth mentioning the closely related line of work involving model reduction for metriplectic systems, which began with Öttinger (2015). As remarked there, preserving metriplecticity in reduced-order models (ROMs) exhibits many challenges due to its delicate requirements on the kernels of the involved operators. There are also hard and soft constrained approaches: the already mentioned Hernández et al. (2021a) aims to learn a close-to-metriplectic data-driven ROM by enforcing degeneracies by penalty, while Gruber et al. (2023a) directly enforces metriplectic structure in projection-based ROMs using exterior algebraic factorizations. The parameterizations of metriplectic data presented here are related to those presented in Gruber et al. (2023a), although NMS does not require access to nonzero components of the gradients $\nabla E,\nabla S$. It is also worth mentioning the influential monograph Dorst et al. (2007) which contains many examples of exterior algebraic methods for computer scientific applications.
@@ 3.1 Exterior algebra
Developing these metriplectic expressions will require some basic facts from exterior algebra, of which more details can be found in, e.g., Tu (2017, Chapter 19). The basic objects in the exterior algebra $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}V$ over the vector space $V$ are multivectors, which are formal linear combinations of totally antisymmetric tensors on $V$. More precisely, if $I(V)$ denotes the two-sided ideal of the free tensor algebra $T(V)$ generated by elements of the form ${}v\otimes~{}v$ (${}v\in V$), then the exterior algebra is the quotient space $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}V\simeq T(V)/I(V)$ equipped with the antisymmetric wedge product operation $\wedge$. This graded algebra is equipped with natural projection operators $P^{k}:\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}V\to\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{k}V$ which map between the full exterior algebra and the $k^{\mathrm{th}}$ exterior power of $V$, denoted $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{k}V$, whose elements are homogeneous $k$-vectors. More generally, given an $n$-manifold $M$ with tangent bundle $TM$, the exterior algebra $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}\mathopen{}\mathclose{{}\left(TM}\right)$ is the algebra of multivector fields whose fiber over $x\in M$ is given by $\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}T_{x}M$.
For the present purposes, it will be useful to develop a correspondence between bivectors $\mathsf{B}\in\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{2}\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right)$ and skew-symmetric matrices ${}B\in\mathrm{Skew}_{n}\mathopen{}\mathclose{{}\left(\mathbb{R}}\right)$, which follows directly from Riesz representation in terms of the usual Euclidean dot product. More precisely, supposing that ${}e_{1},...,~{}e_{n}$ are the standard basis vectors for $\mathbb{R}^{n}$, any bivector $\mathsf{B}\in\mathchoice{{\textstyle\bigwedge}}{{\bigwedge}}{{\textstyle\wedge}}{{\scriptstyle\wedge}}^{2}T\mathbb{R}^{n}$ can be represented as $\mathsf{B}=\sum_{i<j}B^{ij}~{}e_{i}\wedge~{}e_{j}$ where $B^{ij}\in\mathbb{R}$ denote the components of $\mathsf{B}$. Define a grade-lowering action of bivectors on vectors through right contraction (see e.g. Section 3.4 of Dorst et al. (2007)), expressed for any vector ${}v$ and basis bivector ${}e_{i}\wedge~{}e_{j}$ as $\mathopen{}\mathclose{{}\left(~{}e_{i}\wedge~{}e_{j}}\right)\cdot~{}v=\mathopen{}\mathclose{{}\left(~{}e_{j}\cdot~{}v}\right)~{}e_{i}-\mathopen{}\mathclose{{}\left(~{}e_{i}\cdot~{}v}\right)~{}e_{j}.$ It follows that the action of $\mathsf{B}$ is equivalent to